The general inflation rate for oil industry for the 2007 calendar year can be found using the producer price index (from www.bls.gov). When (PPI)2007=267 and Estimated Inflation Rate is 11%, find (PPI)2006?
Answer #1

(PPI2007-PPI2006)/PPI2006=Inflation rate

Hence,
PPI2006=PPI2007/(1+inflation rate)=267/(1+11%)=240.5405405